A tank in the shape of an inverted pyramid has a square base of 4 m on a side and an altitude of 6 m. How many liters of water will it hold when full?
FINAL ANSWER: _______________________________________
1
Expert's answer
2020-12-07T20:26:38-0500

Volume of pyramid = "\frac{1}{3}Ah"


A = base area, h = height


Base area = 4 × 4 = 16m²

Height = 6m


Volume in m³ = "\frac{1}{3}×16×6 = 32m³"

Volume in litres

If 1m³ = 1000litres

Then 32m³ = x litres


=> x = 32 × 1000 = 32 000 litres